If one root of the equation Px+ qx + r = 0 is twice of the other then 2q equals?
A)Pr    B)9Pr   C)P2r   D)Pr2

Grade:12th pass

1 Answers

Vikas TU
6 years ago
Dear Student,
Let one root be m, then the second root is 2m.
Sum of roots= -q/p = m+2m = 3m ------- (1)
Product of roots = r/p = 2m2 ------- (2)
From (2), m2 = r/2p
Squaring (1) and putting this value of m2 in (1),
We get 2q2= 9pr
 
Ans: (C).
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)